# Prompt4Trust: A Reinforcement Learning Prompt Augmentation Framework for Clinically-Aligned Confidence Calibration in Multimodal Large Language Models

### v3

@@ 1 Introduction
Existing confidence calibration techniques attempt to address these challenges through manually designed prompting strategies such as verbalized self-assessments [31], sample consistency measures [14], and reflective prompting [35, 26, 9]. Verbalized confidence methods instruct models to explicitly state their confidence levels, resulting in outputs that align with human expression but still tend to be overconfident [31]. While Xiong et al. [31] unify these approaches into a comprehensive framework employing chain-of-thought prompting, self-random sampling for consistency evaluation, and response aggregation, this requires computationally expensive sampling. Most critically, all existing confidence calibration methods rely on labor-intensive manual prompt engineering to develop prompts and do not explore how the prompts themselves could be directly learned to optimize, not only standard, but clinically informed aspects of calibration.
@@ 2.3 The Limitation of Fixed Prompt Augmentation
A straightforward strategy to approach confidence calibration is to append a fixed auxiliary prompt to the input. For example, one might define $c=$think step-by-step, do not be over-confident", with the expectation that the Downstream Task MLLM $f_{\tau}$ will follow this instruction and produce conservative confidence estimates. However, this heuristic approach suffers from two key limitations: (i) The input text $t\oplus\mathcal{A}$ and fixed auxiliary prompt $c$ are independent; consequently, the auxiliary prompt does not leverage the specific content of $t\oplus\mathcal{A}$. (ii) A fixed prompt often cannot generalize well across distributions of input text $t\oplus\mathcal{A}$ [14, 35].
@@ 3.2 Training CGP Generator Using RL
This reward function encourages $f_{\pi}$ to generate CGPs that guide $f_{\tau}$ toward well-calibrated confidence, while prioritizing the behavior that is most critical to clinical decision-making: when a model expresses high confidence, it should be correct. Correct predictions are rewarded in proportion to their confidence, with higher $\hat{p}$ values yielding higher rewards. Conversely, incorrect predictions are penalized more as confidence increases (via the $1-\hat{p}$ term), and an additional $-1$ penalty is applied. This asymmetry in the reward function differentiates our approach from traditional definitions of calibration and serves two key purposes: (i) it discourages overconfidence errors by assigning the largest penalties to overconfidently incorrect answers, and (ii) it encourages confident expression when the model is correct by reserving the maximal reward for highly confident correct predictions. While some correct predictions may be expressed with lower confidence due to this reward scheme, this behavior is desirable in high stakes settings where conservative uncertainty is preferred over unjustified confidence.
@@ 4.1.2 Dataset
Data splits. We train our model with a randomly selected subset of 5,000 questions from the PMC-VQA training split. Hyperparameter tuning is conducted on a validation set of 1,000 randomly selected samples. Our final results are evaluated on the PMC-VQA-test set which was curated and manually reviewed by Zhang et al. [34, 33].
@@ 4.1.3 Hyperparameters and Training
[…]
@@ 4.1.4 Baselines
  • Verbalized + Fixed-PA: extends the above approach by using the fixed auxiliary prompt augmentation $c=$think step-by-step, do not be over-confident".
@@ 4.1.5 Evaluation Metrics
Method Avg. Confidence on Incorrect Answers $\pm$ std
Verbalized 0.960 ±$\pm 0.135$
Verbalized + Fixed-PA 0.940 ±$\pm 0.155$
Consistency 0.556 ±$\pm 0.301$
Avg-Conf 0.555 ±$\pm 0.301$
Prompt4Trust (Ours) 0.571 ±$\pm 0.180$
@@ Results.
Table 2 reports the average confidence expressed in incorrectly answered responses for Prompt4Trust and all baselines. When incorrect, Prompt4Trust exhibits a low confidence ($0.571$) with low standard deviation (±$\pm 0.180$). In contrast, even when incorrect, the Verbalized and Verbalized + FixedPA baselines express very high confidence ($0.960$ and $0.940$, respectively), which lacks trustworthiness in safety-critical decisions. Furthermore, the low standard deviation in their confidence estimates indicates a consistent pattern of overconfidence. While Consistency and Avg-Conf produce lower average confidence on incorrect predictions ($0.556$ and $0.555$, respectively), their high standard deviations (±$\pm 0.301$) indicate inconsistent and unreliable uncertainty estimates.

### v4

@@ 1 Introduction
footnotetext: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@
Existing confidence calibration techniques attempt to address these challenges through manually designed prompting strategies such as verbalized self-assessments [31], sample consistency measures [14], and reflective prompting [35, 26, 9]. Verbalized confidence methods instruct models to explicitly state their confidence levels, resulting in outputs that align with human expression but still tend to be overconfident [31]. While Xiong et al[31] unify these approaches into a comprehensive framework employing chain-of-thought prompting, self-random sampling for consistency evaluation, and response aggregation, this requires computationally expensive sampling. Most critically, all existing confidence calibration methods rely on labor-intensive manual prompt engineering to develop prompts and do not explore how the prompts themselves could be directly learned to optimize, not only standard, but clinically informed aspects of calibration.
@@ 2.3 The Limitation of Fixed Prompt Augmentation
A straightforward strategy to approach confidence calibration is to append a fixed auxiliary prompt to the input. For example, one might define $c=$ "think step-by-step, do not be over-confident", with the expectation that the Downstream Task MLLM $f_{\tau}$ will follow this instruction and produce conservative confidence estimates. However, this heuristic approach suffers from two key limitations: (i) The input text $t\oplus\mathcal{A}$ and fixed auxiliary prompt $c$ are independent; consequently, the auxiliary prompt does not leverage the specific content of $t\oplus\mathcal{A}$. (ii) A fixed prompt often cannot generalize well across distributions of input text $t\oplus\mathcal{A}$ [14, 35].
@@ 3.2 Training CGP Generator Using RL
This reward function encourages $f_{\pi}$ to generate CGPs that guide $f_{\tau}$ toward well-calibrated confidence, while prioritizing the behavior that is most critical to clinical decision-making: when a model expresses high confidence, it should be correct. Correct predictions are rewarded in proportion to their confidence, with higher $\hat{p}$ values yielding higher rewards. Conversely, incorrect predictions are penalized more as confidence increases (via the $1-\hat{p}$ term), and an additional 1-1 penalty is applied. This asymmetry in the reward function differentiates our approach from traditional definitions of calibration and serves two key purposes: (i) it discourages overconfidence errors by assigning the largest penalties to overconfidently incorrect answers, and (ii) it encourages confident expression when the model is correct by reserving the maximal reward for highly confident correct predictions. While some correct predictions may be expressed with lower confidence due to this reward scheme, this behavior is desirable in high stakes settings where conservative uncertainty is preferred over unjustified confidence.
@@ 4.1.2 Dataset
Data splits. We train our model with a randomly selected subset of 5,000 questions from the PMC-VQA training split. Hyperparameter tuning is conducted on a validation set of 1,000 randomly selected samples. Our final results are evaluated on the PMC-VQA-test set which was curated and manually reviewed by Zhang et al[34, 33].
@@ 4.1.3 Hyperparameters and Training
A grid search over temperature, top-k, and beta is conducted to obtain optimal hyperparameter configuration. We report all results using the optimal hyperparameter configuration ($\texttt{beta}=0.04,\texttt{top-k}=100,\texttt{temperature}=0.6$) which favorably balances calibration and accuracy.
@@ 4.1.4 Baselines
  • Verbalized + Fixed-PA: extends the above approach by using the fixed auxiliary prompt augmentation $c=$ "think step-by-step, do not be over-confident".
@@ 4.1.5 Evaluation Metrics
Method
Avg. Confidence on
Incorrect Answers $\pm$ std
Verbalized 0.960 $\pm 0.135$
Verbalized + Fixed-PA 0.940 $\pm 0.155$
Consistency 0.556 $\pm 0.301$
Avg-Conf 0.555 $\pm 0.301$
Prompt4Trust (Ours) 0.571 $\pm 0.180$
@@ Results.
Table 2 reports the average confidence expressed in incorrectly answered responses for Prompt4Trust and all baselines. When incorrect, Prompt4Trust exhibits a low confidence ($0.571$) with low standard deviation ($\pm 0.180$). In contrast, even when incorrect, the Verbalized and Verbalized + FixedPA baselines express very high confidence ($0.960$ and $0.940$, respectively), which lacks trustworthiness in safety-critical decisions. Furthermore, the low standard deviation in their confidence estimates indicates a consistent pattern of overconfidence. While Consistency and Avg-Conf produce lower average confidence on incorrect predictions ($0.556$ and $0.555$, respectively), their high standard deviations ($\pm 0.301$) indicate inconsistent and unreliable uncertainty estimates.